OpenStudy (anonymous):

lim tan3x/ 3tan2x x->0

OpenStudy (anonymous):

Is it zero? Separate the two and multiply out. It is either zero or indeterminate, but I'd go with zero.

OpenStudy (anonymous):

0.

OpenStudy (anonymous):

the answerr is 1/2

OpenStudy (anonymous):

use l' hospital rule and u'll get 1/2

OpenStudy (anonymous):

Just checked it on Wolfram Alpha. The answer is 0.

OpenStudy (anonymous):

Using the l'Hôpital's rule: Take derivative of the top and the bottom: tan3x = 3sec²3x 3tan2x = 6sec²2x So its: 3sec²3x/6sec²2x Now plug in x=0 You'll get 3(1)/6(1) = 3/6 =1/2

OpenStudy (anonymous):

Thanks all for your help. Any solution without that rule for now? Because the teacher said that we should be able to do this question without any rules. I wasnt taught l'hopitals rule yet

OpenStudy (anonymous):

then using trigonometic formulaes u can xpand it and then u'll get what u want...:) all d best..:)

OpenStudy (anonymous):

@bobobobobb so then i'm gonna have to say, go with zero

OpenStudy (anonymous):

It would be zero. Divide and conquer; veni vidi vici. Tan... * tan... will both go to zero, multiply out and you will get zero.

OpenStudy (anonymous):

no i just solved it with trigonometric formulaes i'm getting 1/2

OpenStudy (anonymous):

the answer key says 1/2 but i dont know how

OpenStudy (anonymous):

see in numerator u xpand tan 3x and do the same for 6tan 2x in denominator and then put x=0 and u'll get 1/2

OpenStudy (anonymous):

expand? what does that mean?

OpenStudy (anonymous):

means apply formula for tan 3x and u'll get everything in tan x same u do in denominator and then put x=0 to get ur result

OpenStudy (anonymous):

okay. thank you very much